At the beginning of Year 1, Oak Consulting had the following normal balances in its accounts:

Account Balance
Cash $ 32,200
Accounts receivable 16,600
Accounts payable 10,500
Common stock 27,400
Retained earnings 10,900


The following events apply to Oak Consulting for Year 1:

  1. Provided $67,000 of services on account.
  2. Incurred $2,900 of operating expenses on account.
  3. Collected $48,400 of accounts receivable.
  4. Paid $37,500 cash for salaries expense.
  5. Paid $12,060 cash as a partial payment on accounts payable.
  6. Paid a $10,000 cash dividend to the stockholders.

c. Show the beginning balances and the events in a horizontal statements model such as the following one: (In the Statement of Cash Flows column, use the initials "OA" for operating activities, "FA" for financing activities and "NC" for net change in cash. Select "NA" wherever required. Enter any decreases to account balance and cash outflows with a minus sign.)

OAK CONSULTING
Effect of Transactions on the Financial Statements for Year 1
Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders' Equity Revenue Expenses = Net Income
No Cash + Accounts Receivable = Accounts Payable + Common Stock + Retained Earnings
Bal. + = + + =
1 + = + + =
2 + = + + =
3 + = + + =
4 + = + + =
5 + = + + =
6 + = + + =
Bal. 0 + 0 = 0 + 0 + 0 0 0 = 0 0
Answer #1

c. Show the beginning balances and the events in a horizontal statements model such as the following one: (In the Statement of Cash Flows column, use the initials "OA" for operating activities, "FA" for financing activities and "NC" for net change in cash. Select "NA" wherever required. Enter any decreases to account balance and cash outflows with a minus sign.)

OAK CONSULTING
Effect of Transactions on the Financial Statements for Year 1
Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders' Equity Revenue Expenses = Net Income
No Cash + Accounts Receivable = Accounts Payable + Common Stock + Retained Earnings
Bal. 32200 + 16600 = 10500 + 27400 + 10900 =
1 + 67000 = + + 67000 67000 = 67000
2 + = 2900 + + -2900 -2900 = -2900
3 48400 + -48400 = + + = 48400 OA
4 -37500 + = + + -37500 -37500 = -37500 -37500 OA
5 -12060 + = -12060 + + = -12060 OA
6 -10000 + = + + -10000 = -10000 FA
Bal. 21040 + 35200 = 1340 + 27400 + 27500 67000 40400 = 26600 -11160